OpenStudy (katielong):

x = 1/6 f(2x) can be expressed in the form a+bx find the values of a and b

jhonyy9 (jhonyy9):

if f(x)=1/6 than f(2x)=2(1/6)=2/6=1/3

OpenStudy (katielong):

is that right?